OpenStudy (anonymous):

Add: (8x + 2y – 3z – 4) + (–4x – 5y + 4z + 6)

OpenStudy (anonymous):

please helpp

OpenStudy (anonymous):

@navk please help

OpenStudy (anonymous):

First combine the like terms and then add them. do you know what the like terms are

OpenStudy (anonymous):

yeah,i know what they are @navk

OpenStudy (anonymous):

so for example you would first combine the x terms to get 8x - 4x = 4x then combine the y , z and constant terms similarly

OpenStudy (anonymous):

so would it be 7y and 7x? @navk

OpenStudy (anonymous):

no how did you get that can you show your work

OpenStudy (anonymous):

i just added them together

OpenStudy (anonymous):

2y-5y= -3y right and then 3z +4z=7z @navk

OpenStudy (anonymous):

Great! That is correct except that the 3x has a negative sign on it.

OpenStudy (anonymous):

so do i take off the negative sign or does it stay there? @navk

OpenStudy (anonymous):

You can't change the question itself! Or are you talking about the answer, because in it there is no negative sign to take off! You need to put a negative sign before the 3z

OpenStudy (neer2890):

it's 8x-4x+2y-5y-3z+4z-4+6 =4x-3y+z+2

OpenStudy (anonymous):

thanks @neer2890 even though i didnt want you to work it out for me,i wanted to figure it out on my own mostly

OpenStudy (neer2890):

ok then try some other equation buddy